OpenStudy (anonymous):

find the solution of the equation x^2-12x+117 I keep getting this damn thing wrong

OpenStudy (nurali):

http://www.geteasysolution.com/x%5E2+12x+117=0

OpenStudy (anonymous):

The equation has no real roots. The curve does not touche or cross the x axis. Refer to the attached plot.

OpenStudy (anonymous):

but no solution is not part of the multiple choices...

OpenStudy (agreene):

As stated above, there are no real roots. \[x=6\pm9i\] are the complex roots.

OpenStudy (anonymous):

alright, not sure how that you got to the solution.

OpenStudy (anonymous):

x^2-12x+117 solution : deta = b²-4ac deta= (-12)²-4*1*117 deta= -324 = 18i x= 12+-18i/2 = 6+9i or 6-9i

OpenStudy (agreene):

\[x^2-12 x+117 = 0\] \[x^2-12 x = -117\] \[x^2-12 x+36 = -81\] \[(x-6)^2 = -81\] \[x-6 = 9 i\;\;\text{or}\;\;\;x-6 = -9 i\] \[x = 6+9 i\;\;\;\text{ or }\;\;\;\ x = -9 i+6\] written another way...\(x= 6\pm9i\)

OpenStudy (anonymous):

oh, okay, I think I'm getting a hang of it